In the label, it is written that a light bulb has a mean life of 360 and with a standard deviation of 90 hours. To test the claim, 60 light bulbs were tested and showed a mean life of 354. At the 0.05 significance level, can we say that mean life is different from 360?


1
Expert's answer
2021-07-15T09:37:58-0400

"\mu=360 \\

\sigma=90 \\

n=60 \\

\bar{x}=354 \\

H_0: \mu=360 \\

H_1: \mu ≠360"

Test-statistic:

"Z = \frac{\bar{x} - \mu}{\sigma/ \sqrt{n} } \\

Z = \frac{354 - 360}{90 / \sqrt{60}}= -0.516 \\

α=0.05"

Two-tailed test

"Z_{crit}=1.96"

The decision rule is: Reject H0 if Z ≤ -1.96 or if Z ≥ 1.96.

"Z= -0.516 > Z_{crit} = -1.96"

Accept H0.

There is enough evidence to conclude that light bulb has a mean life of 360 at the 0.05 significance level.